OpenStudy (jkbo):

Please help (x · y) · z = (y · x) · z is an example of which property?  A.Associative of Multiplication  B.Identity of Multiplication  C.Distributive  D.Commutative of Multiplication 

OpenStudy (anonymous):

Associative of multiplication is a(b x c)=c(a x b) Identity of multiplication is a x 1=a Distributive property is a(b+c)=ab+ac Commutative of multiplication is a x b=b x a I hope that helps

OpenStudy (jkbo):

ok so its A @chrffl

OpenStudy (anonymous):

yes

Directrix (directrix):

@jkbo @chrffl A. Associative of Multiplication is NOT correct.

OpenStudy (anonymous):

Would it be commutative then?

Directrix (directrix):

The Associative Property is the grouping property. Commutative Property has to do with order. If you look at the problem, (x · y) · z = (y · x) · z, the grouping of x and y together has not changed. So, the answer is not Association. The order of the x and y has changed. So, yes, Commutative Property. @chrffl

OpenStudy (anonymous):

Wow, I can't believe I missed that. Thanks for correcting me